OpenStudy (anonymous):

question on dimensional analysis, 1/2mv^2=1/2mv0^2+sq root of mgh need to determine why cant be correct no value given for h, i assume m is mass, also g is gravity any idea for h

OpenStudy (abb0t):

\(\frac{1}{2}mv^2=\frac{1}{2}mv_0^2+ \sqrt{mgh}\) all you are doing here is solving for your variable \(h\), length.

OpenStudy (akashdeepdeb):

H is height and the only reason the second term on the RHS cannot be square rooted is because when its dimensional formula is found it does not match with the other 2! Understood? :)